# Bone: Block-Affine Adaptation of Large Language Models

### v5

@@ 4.2 Experiment on Natural Language Generation
The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Bone demonstrates outstanding performance across three different tasks. On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants. On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
@@ 4.4 Bone vs Bat
To validate the effectiveness of the “Weight Guide”, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K. As shown in the table, Bat, equipped with “Weight Guide”, achieves significant improvements in performance metrics compared to Bone.
@@ 4.5 Ablation experiments on different grouping methods for Bat
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bat fine-tuning. This allows for a fairer comparison between Bat-row and Bat-col. The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
@@
4.6 Ablation experiments on the specific implementation of “Weight Guide”
Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
The formula for Bone-Hadamard is as follows:
$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (5)
$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\odot bone_{d/b,b,b}+bone_{d/b,b,b})$
Table 6: Block-Affine on math tasks
Model Strategy Trainable GSM8K Math
RWKV6-3B Bat 55.1M 25.93 3.32
Bat-Hadamard 55.1M 22.44 2.67
@@ 5 Resource and efficiency
Table  7 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
@@
Therefore, a key focus of our future work will be improving the Bat operator to enhance token throughput and reduce memory usage.
@@ 6 Conclusion
This work, inspired by GQA and MQA, leverages the sparsity of LLM weights to design the Block-Affine Adaptation (Bone) structure. In Bone, the original weights are divided into multiple subspaces, all of which share a single low-rank matrix initialized to zero for updates. Extensive experiments demonstrate that Bone consistently outperforms LoRA and its variants across various tasks, while also offering superior computational efficiency. To break the limitations of LoRA and effectively utilize the information from the original weights, we propose the “Weight Guide” theory, which enables significant improvements through simple linear transformations.
By integrating these innovations, we introduce a new structure called Block-Affine Transformation (Bat). Bat not only validates the effectiveness of “Weight Guide” but also addresses the limitation of Bone, where identical updates are applied to all subspaces. Experimental results show that Bat surpasses Bone on multiple tasks, although with a trade-off in computational efficiency.
Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
@@ 7 Future work
Bone and “Weight Guide” are merely a starting point, offering a foundation for researchers to explore additional branches inspired by Bat. There are many directions for future work.
1. Can Bone adapt to multimodal tasks and other complex scenarios?
2. Can “Weight Guide” be implemented using alternative computation methods?

### v6

@@ 4.2 Experiment on Natural Language Generation
Figure 5: The image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
@@
The results, as shown in Table  2 and Figure  5, demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Bone demonstrates outstanding performance across three different tasks. On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants. On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
@@ 4.4 Bone vs Bat
In this section, we have verified the effectiveness of non-linear updates by comparing the performance of Bone and Bat, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K. As shown in the table, with the same number of trainable parameters, the metrics of non-linear updates are better, successfully breaking the original limits.
@@ 4.5 Ablation experiments on different grouping methods for Bat
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bat fine-tuning. This allows for a fairer comparison between Bat-row and Bat-col.
@@
The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
@@ 5 Resource and efficiency
Table  6 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
@@ 6 Conclusion
In this work, we proposed the DiSHA framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. Building on the DiSHA framework, we introduced Bone, a broad and efficient method comparable to LoRA. Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
To address the potential issue of linear correlation in shard updates within Bone and the DiSHA framework, we proposed Bat. By applying a linear transformation between the trainable matrix and the pre-trained weights, Bat ensures nonlinear independence in the updates of different shards, successfully breaking through the structural limitations of Bone
@@ 7 Future work
Due to the limitations of the experimental equipment, we did not report any results from the full-scale training required for this work. However, through extensive evaluations and comparisons with LoRA and its variants, we validated the effectiveness of DiSHA, Bone, and Bat. The outstanding performance demonstrated in LLMs by this work gives us great confidence. We believe that DiSHA, Bone, and Bat can be applied to any multimodal task. Of course, this work also has some limitations that need to be addressed. Although Bat successfully overcomes the limitations of Bone, it also leads to reduced computational efficiency and increased resource consumption. Therefore, in the future, we plan to develop specialized optimized operators to address this issue.
@@
8 Acknowledgments
I would like to express my heartfelt gratitude to all the seniors who provided valuable feedback on my work. Their insightful suggestions have greatly enhanced the quality of this paper. I also wish to thank RWKVOS for their support. Finally, I sincerely appreciate the meticulous guidance provided by the anonymous online mentor, Professor Cormen, whose help has been indispensable.